OpenStudy (anonymous):

Determine the area under the standard normal curve that lies to the left of the following z value: z = –0.33 Use a table if necessary. Round your answer to the nearest ten thousandth.

OpenStudy (anonymous):

Please help me... :(

OpenStudy (amistre64):

what do you have to use? calculator, program, table, etc ...

OpenStudy (amistre64):

integrating the function?

OpenStudy (amistre64):

\[\Large \frac{1}{\sigma\sqrt{2\pi}}~\int_{a}^{b}e^{\frac{(x-\mu)^2}{2\sigma^2}}dx\] and since a normal curve without any other information is assumed to have a u=0 and sd=1 \[\Large \frac{1}{\sqrt{2\pi}}~\int_{a}^{b}e^{\frac{x^2}{2}}dx\]

OpenStudy (anonymous):

can you go through the steps? Sorry, but I am so confused

OpenStudy (anonymous):

I don't have a calculator, program or table :(

OpenStudy (amistre64):

i need to know what you have to do this with; ideally a table or a ti83

OpenStudy (anonymous):

we need to find it on a table I believe

OpenStudy (amistre64):

hmm, a table can be found online, goole ztable

OpenStudy (amistre64):

http://statstutorstl.blogspot.com/2010/07/z-table-gives-probabilty-distribution.html heres one thatll do fine

OpenStudy (anonymous):

0.3707?

OpenStudy (amistre64):

maybe, how did you get to that?

OpenStudy (anonymous):

I went to -0.33 and used the number provided

OpenStudy (amistre64):

there was a -0.33 ? onthe table i linked to; they measure from the mean; so if we cross hairs 0.3 with .03 we get .6293, which is .5000 to big

OpenStudy (anonymous):

you use both the colum on top and on the side

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

I didn't use the table you posted

OpenStudy (amistre64):

which table did you use?

OpenStudy (anonymous):

you didn't look on the negative portion of the chart

OpenStudy (anonymous):

heres another question Determine the area under the standard normal curve that lies to the left of the following z score. Use a table if necessary. Round your answer to four decimal places. z = 0.33

OpenStudy (amistre64):

ah, i see my error;